OpenStudy (anonymous):

graphing form : containing ( 2, -5) and ( 4, 4)

OpenStudy (anonymous):

find the slope (-5-4)/(2-4) = -9/-2 = 9/2 put it into point slope form: y-4 = 9/2 (x-4) you can graph from that form. or solve for y. y-4=9/2x-18 .................add 4 y=9/2x-14

OpenStudy (anonymous):

didnt we just answer this one...?

OpenStudy (anonymous):

my paper says to find the standard form , graphing form and , point slope so idk if i only need one answer and do the same work for the rest of them?

OpenStudy (amistre64):

graphing form?

OpenStudy (anonymous):

graphing form must be slope intercept form, but you can graph in all 3 forms if you know how.

OpenStudy (anonymous):

you can get all forms from the same equation, any different form us just a different way of writing the same equation

OpenStudy (anonymous):

take y=9/2x+14. it's in slope intercept (graphing) form.

OpenStudy (anonymous):

get x and y on the same side by subtracting 9/2x from both sides. -9/2x+y=14. you can't have fractions or a (-) x. so multiply equation by -2 9x-2y=-28..........................standard.

OpenStudy (anonymous):

okay so if i decide to graph them would i put 9/2 as rise over run or (x-4)?